# Explainable Molecular Property Prediction: Aligning Chemical Concepts with Predictions via Language Models

### v3

@@ 3.2 Theoretical Analysis
Manifold Hypothesis. It is widely believed that natural data, including molecules, distribute around a manifold (Bordt et al., 2023; Lin et al., 2022; Godwin et al., 2022; Singh et al., 2020). According to the manifold hypothesis for gradient-based explanations (Bordt et al., 2023), if a feature lies in the tangent space of a manifold, then the feature respects the manifold and contributes to the class, and such a feature is desirable to be explained. We call these features "causal features" in our work. Conversely, if a feature is orthogonal to the manifold, then the feature does not contribute to the class. We call these features "spurious features".
@@
With this decomposition, we demonstrate how the gradient-based explanations $\nabla_{g}\log p(y|g)$ can uncover the structure-property relationships. Due to the page limitation, we provide the proof in Appendix.
@@
4 Experiments
4.1 Experimental Setup
Datasets. We use six datasets on two types of tasks, i.e., hepatotoxicity and mutagenicity, to evaluate the algorithmic performance for the explainable molecular property prediction. Six mutagenicity datasets are Mutag (Debnath et al., 1991), Mutagen (Morris et al., 2020), PTC-FM (Toivonen et al., 2003), PTC-FR (Toivonen et al., 2003), PTC-MM (Toivonen et al., 2003), and PTC-MR (Toivonen et al., 2003). For hepatotoxicity (Toivonen et al., 2003), the Liver dataset (Liu et al., 2015) is used. We used these datasets as the ground truth substructures in these datasets are known. For detailed information on these datasets and the ground truth substructures, please refer to Appendix.
@@ 4.1 Experimental Setup
Evaluation Metrics. We use three metrics for evaluating the performance of the proposed Lamole. 1) Classification Accuracy: We evaluate the model’s predictions by $\sum_{i=1}^{I}\mathbb{I}(y^{(i)}=\hat{y}^{(i)})/I$. 2) Explanation Accuracy: We follow the experimental settings in GNNExplainer (Ying et al., 2019), which formulates the explanation problem as a binary classification of edges. We treat edges inside ground-truth substructure as positive edges and negative otherwise, and AUC is adopted as the metric for quantitative evaluation. We only consider the mutagenic/hepatotoxic molecules because no explicit substructures exist in nonmutagenic/nonhepatotoxic ones. 3) Explanations’ Plausibility: We use the defined explanations’ plausibility $\mathrm{EP}$ to measure how confident the explanation aligns with the ground truth.
@@
Baselines. We combine Lamole into three BERT family models, such as DistilBert (Sanh, 2019), DeBerta (He et al., 2020), and Bert to evaluate the performance of the proposed Lamole. For evaluating classification accuracy, we compare our Lamole with one SMILES string-based LM, ChemBERTa (Chithrananda et al., 2020) and several GNNs including GCN (Kipf & Welling, 2016), DGCNN (Zhang et al., 2018), edGNN (Jaume et al., 2019), GIN (Xu et al., 2018), RW-GNN (Nikolentzos & Vazirgiannis, 2020), DropGNN (Papp et al., 2021), and IEGN (Maron et al., 2018).
@@
For evaluating explanation accuracy, Lamole is compared with three types of alternative methods: 1) GCN with feature-based explainability techniques, including SmoothGrad (Smilkov et al., 2017), GradInput (Shrikumar et al., 2017), and GradCAM (Selvaraju et al., 2017), 2) Bert with the above feature-based explainability techniques, where Group SELFIES is used as input for a fair comparison, and 3) explainable GNNs including OrphicX (Lin et al., 2022), GNNExplainer (Ying et al., 2019), PGExplainer (Luo et al., 2020), and Gem (Lin et al., 2021). The details of experimental settings can be found in Appendix.
4.2 Results
Prediction Performance. Table 1 shows the classification accuracy of compared algorithms. As we can see, our proposed Lamole+DistilBert, Lamole+DeBerta, and Lamole+Bert not only can provide explainability but also can achieve comparable prediction accuracy as compared to existing predictive methods. In addition, Lamole models show superior performance over ChemBERTa. This suggests that using molecular representations with more chemical semantics, like Group SELFIES, can help LMs better learn the chemical semantics and structure-property relationships.
Table 1: Mean Classification Accuracy on the Seven Datasets (%)
Methods Mutag Mutagen PTC-FM PTC-FR PTC-MM PTC-MR Liver
GCN (Kipf & Welling, 2016) 84.6 78.9 54.8 63.0 57.8 53.3 41.1
DGCNN (Zhang et al., 2018) 85.8 74.8 57.3 63.5 61.0 58.6 44.6
edGNN (Jaume et al., 2019) 86.9 75.2 59.8 65.7 64.4 56.3 44.5
GIN (Xu et al., 2018) 87.5 82.3 62.1 66.2 65.1 64.0 44.9
RW-GNN (Nikolentzos & Vazirgiannis, 2020) 87.2 80.3 61.9 64.0 62.4 57.0 43.2
DropGNN (Papp et al., 2021) 89.4 80.7 62.0 66.0 63.7 64.2 45.0
IEGN (Maron et al., 2018) 84.6 80.1 60.8 59.8 61.1 59.5 45.3
ChemBERTa (Chithrananda et al., 2020) 86.8 78.0 60.0 65.7 60.4 58.7 45.7
@@ 4.2 Results
Explanation Performance. Table 2 presents the explanation accuracy of the compared explainability techniques. It should be noted that the ground truth annotations used in our work provide additional supervisory signals. Therefore, we also align the generated explanations of these baselines with the annotations for fair comparison, and the ground truth annotation rate is 10%. Lamole improves the explanation accuracy by $1.4\%\sim 14.3\%$ compared to the baseline methods. We also investigated the explanation accuracy of our Lamole under different ground truth annotation rates (10%, 20%, 50%, and 100%). The impact of annotation rates can be found in Appendix. In addition, we discuss the rationale of using labeled annotations (see Appendix).
Table 2: Mean Explanation Accuracy on the Seven Datasets (%)
Methods Mutag Mutagen PTC-FM PTC-FR PTC-MM PTC-MR Liver
GradInput+GCN (Shrikumar et al., 2017) 70.3 67.9 69.7 66.4 64.6 65.0 73.0
GradCAM+GCN (Selvaraju et al., 2017) 69.8 67.0 71.0 67.9 66.2 67.3 69.4
SmoothGrad+GCN (Smilkov et al., 2017) 69.2 66.8 67.5 62.6 64.9 63.1 66.4
GradInput+Bert (Shrikumar et al., 2017) 75.1 72.6 73.0 68.9 65.6 69.6 75.6
GradCAM+Bert (Selvaraju et al., 2017) 75.3 72.4 77.5 70.0 70.2 73.0 76.0
SmoothGrad+Bert (Smilkov et al., 2017) 73.4 72.8 73.7 71.0 67.0 69.9 75.1
GNNExplainer (Ying et al., 2019) 70.6 64.2 68.9 67.9 66.8 67.1 72.1
PGExplainer (Luo et al., 2020) 66.5 58.7 70.3 68.0 65.9 67.0 71.5
Gem (Lin et al., 2021) 73.7 66.0 71.3 69.0 68.9 69.2 73.6
OrphicX (Lin et al., 2022) 78.0 71.5 74.6 70.4 70.9 71.4 74.0
@@
Figure 4: Explanation visualization of one molecule (ID: 155) from the Mutag dataset.
@@
We selected some representative molecules for explanation visualization. These explanations are shown in Figs. 1 and 4, respectively. The right panel of those figures is the importance scores obtained by Lamole, where "other" in the figures is the average importance score of other unlisted functional groups/fragments. Compared to baseline methods, Lamole provides chemically meaningful explanations. Particularly, the interaction among the functional groups is successfully captured. More visualization results can be found in Appendix.
In addition, we evaluated the performance of compared algorithms by using the proposed explanation plausibility metric $\mathrm{EP}$. The statistical results of $\mathrm{EP}$ are presented in Figs. 5 and LABEL:epfr. From the figures, we can observe that the $\mathrm{EP}$ values of the comparison algorithm are slightly lower, which means that the algorithms cannot confidently reflect the relationships between structure and property. Compared to the comparison algorithm, the $\mathrm{EP}$ values of Lamole have increased by $2\%\sim 9\%$. More analysis regarding the explanation plausibility can be found in Appendix.
4.3 Ablation Studies
We conducted ablation studies for each component in our Lamole. Specifically, we removed the attention weights in the explanations, removed the marginal loss, and only used attention weights as explanations. The corresponding ablation algorithms are named Lamole ($-\alpha$), Lamole ($-\mathcal{L}_{M}$), and Lamole (Att), respectively. The results are shown in Fig. 6. The explanation accuracy decreases by 1.4$\sim$2.3when removing the attention weights. Removing the marginal loss can decrease the explanation accuracy by 1.0$\sim$5.0%. Regarding the results of using only attention weights, The explanation accuracy decreases by 1.4$\sim$6.3%. The above results confirm the effectiveness of using the marginal loss, attention weights, and gradients.
Figure 6: The explanation accuracy of Lamole+Bert, Lamole+Bert ($-\alpha$), Lamole+Bert ($-\mathcal{L}_{M}$), and Lamole (Att).
From the perspective of model training, the marginal loss enables the model to be trained under the causal signals. To verify, we compared the classification performance with and without the marginal loss, as shown in Table 3. Without the marginal loss, the classification accuracy degrades by 0.7$\sim$3.7%. The above results indicate that marginal loss could help identify the causal features, thereby improving classification accuracy.

### v4

@@ III-B Theoretical Analysis
Manifold Hypothesis. It is widely believed that natural data, including molecules, distribute around a manifold [5, 30, 14, 47]. According to the manifold hypothesis for gradient-based explanations [5], if a feature lies in the tangent space of a manifold, then the feature respects the manifold and contributes to the class, and such a feature is desirable to be explained. We call these features "causal features" in our work. Conversely, if a feature is orthogonal to the manifold, then the feature does not contribute to the class. We call these features "spurious features".
@@
With this decomposition, we demonstrate how the gradient-based explanations $\nabla_{g}\log p(y|g)$ can uncover the structure-property relationships.
@@
Proof of Theorem 1: The gradient with respect to the prediction $\nabla_{g}\log p(y|g)$ can be decomposed into the gradient on the causal features and spurious features, respectively,
$\nabla_{g}\log p(y|g)=\nabla_{g}\log p(y|g)\odot\mathbf{m}(g)\\ +\nabla_{g}\log p(y|g)\odot(1-\mathbf{m}(g)).$ (7)
By minimizing the loss of Eq. (5), the gradients on spurious features $\nabla_{g}\log p(y|g)\odot(1-\mathbf{m}(g))$ are suppressed, and $\nabla_{g}\log p(y|g)$ approximates $\nabla_{g}\log p(y|g)\odot\mathbf{m}(g)$. Therefore, we have $\nabla_{g}\log p(y|g)\approx\nabla_{g}\log p(y|g)\odot\mathbf{m}(g)$. On the other hand, $\nabla_{g}\log p(y|g)\odot\mathbf{m}(g)$ can be rewritten as
$\nabla_{g}\log p(y|g)\odot\mathbf{m}(g)=\nabla_{g}\log p(g|y)\odot\mathbf{m}(g)\\ -\sum_{j}p(y=j|g)\nabla_{g}\log p(g|y=j)\odot\mathbf{m}(g).$ (8)
Because the data distribution $p(g|y)\odot\mathbf{m}(g)$ reflects the causal feature manifold $\mathcal{M}_{c}$, the gradient of the distribution $\nabla_{g}p(g|y)\odot\mathbf{m}(g)$ represents the tangent space of the causal feature manifold $\mathcal{M}_{c}$. In addition, Eq. (8) shows that the $\nabla_{g}\log p(y|g)\odot\mathbf{m}(g)$ is a linear combination of $\nabla_{g}p(g|y)\odot\mathbf{m}(g)$, so $\nabla_{g}\log p(y|g)\odot\mathbf{m}(g)$ also lies tangent space of the manifold $\mathcal{M}_{c}$.
Together with Eq. (7) and Eq. (8), we prove the gradient-based explanations $\nabla_{g}\log p(y|g)$ lies tangent space of the manifold $\mathcal{M}_{c}$. This indicates by minimizing the loss of Eq. (5), the model $p(y|g)$ has reflected the causal feature manifold. According to the manifold hypothesis, the features on the causal feature manifold contribute to the molecular property. Therefore, the gradient-based explanations $\nabla_{g}\log p(y|g)$ can uncover the causal features, thus revealing the structure-property relationships. This completes the proof.
IV Experiments
IV-A Experimental Setup
Datasets. We use six datasets on two types of tasks, i.e., hepatotoxicity and mutagenicity, to evaluate the algorithmic performance for the explainable molecular property prediction. Six mutagenicity datasets are Mutag [9], Mutagen [35], PTC-FM [50], PTC-FR [50], PTC-MM [50], and PTC-MR [50]. For hepatotoxicity [50], the Liver dataset [31] is used. Larger-sized molecules typically include more complex structures. The datasets that we used contained relatively large molecules. The maximal number of atoms of Mutag, Mutagen, PTC-FM, PTC-FR, PTC-MM, PTC-MR, and Liver are 26, 417, 64, 64, 64, 64, and 157, respectively. The details of the used dataset are provided in Table I.
TABLE I: Statistiscal Information of the Datasets
Datasets Mutag Mutagen PTC-FM PTC-FR PTC-$M$ PTC-MR Liver
Graphs 188 4337 349 351 336 344 587
Classes 2 2 2 2 2 2 3
Max nodes 26 417 64 64 64 64 157
Avg nodes 17.9 29 14.1 14.6 14 14.3 25.6
Avg edges 19.8 30 14.5 15 14.3 14.7 27.4
Ground truth 120 724 58 49 51 61 187
  • denotes the number of molecules with known ground truth substructures.
Following OrphicX [30], on the Mutag, Mutagen, PTC-FM, PTC-FR, PTC-MM, and PTC-MR datasets, we only consider the explanations for the mutagenic class, because the molecules of the non-mutagenic class have no ground truth. Although some works used single $\mathrm{N}=\mathrm{N}$, $\mathrm{NO_{2}}$, or $\mathrm{NH_{2}}$ as ground truth, this is not reasonable, as 32% of non-mutagenic graphs in Mutagen containing at least single $\mathrm{NO_{2}}$ or $\mathrm{NH_{2}}$. In fact, the ground truth for the mutagenic class is the benzene with a chemical group on it, such as $\mathrm{N}=\mathrm{N}$, $\mathrm{NO_{2}}$, and $\mathrm{NH_{2}}$ [29, 30, 38].
For the Liver dataset, the molecules of possible hepatotoxicity with ground truth substructures and hepatotoxicity with ground truth substructures are collected for explainable molecular property prediction. The twelve ground truth substructures of the Liver dataset are shown in Fig. 4.
To assess the generalizability of our method’s explainable prediction ability, we performed an extended evaluation on two critical absorption, distribution, metabolism, and excretion (ADME) properties: solubility and permeability. The evaluation was conducted using the ZINC dataset [18]. Lacking expert-annotated references for these complex endpoints, we employed an automated, descriptor-driven approach to generate the required ground-truth functional group labels. The annotation procedure for the ground truth substructures in the dataset is detailed in the supplemental material.
Figure 4: Twelve ground truth substructures of the Liver dataset. Lowercase element symbols represent aromatic atoms of the element; the letter "a" matches any aromatic atom. Elements in square brackets match any of the elements in a molecule.
@@ IV-A Experimental Setup
Evaluation Metrics. We use four metrics for evaluating the performance of the proposed Lamole. 1) Classification Accuracy: We evaluate the model’s predictions by $\sum_{i=1}^{I}\mathbb{I}(y^{(i)}=\hat{y}^{(i)})/I$. 2) Explanation Accuracy: We follow the experimental settings in GNNExplainer [64], which formulates the explanation problem as a binary classification of edges. We treat edges inside ground-truth substructure as positive edges and negative otherwise, and AUC is adopted as the metric for quantitative evaluation. We only consider the mutagenic/hepatotoxic molecules because no explicit substructures exist in nonmutagenic/nonhepatotoxic ones. 3) Explanations’ Fidelity: If an explanation is faithful, removing the most important fragments identified can cause a significant change in the model’s prediction score for the target property. The definition is as follows: $\text{Fidelity}(G,G_{\text{ex}})=[f(G)]_{c^{*}}-[f(G/G_{\text{ex}})]_{c^{*}}$, where $G$ is the original molecular inputs, $G_{\text{ex}}$ is the set of key functional groups/fragments identified, and $f(\cdot)$ is the model’s classification probability or regression output. 4) Explanations’ Plausibility: We use the defined explanations’ plausibility $\mathrm{EP}$ to measure how confident the explanation aligns with the ground truth.
@@
Baselines. We combine Lamole into three BERT family models, such as DistilBert [42], DeBerta [15], and Bert to evaluate the performance of the proposed Lamole. For evaluating classification accuracy, we compare our Lamole with one SMILES string-based LM, ChemBERTa [7] and several GNNs including GCN [23], DGCNN [66], edGNN [21], GIN [62], RW-GNN [36], DropGNN [37], and IEGN [34].
@@
For evaluating explanation accuracy, Lamole is compared with three types of alternative methods: 1) GCN with feature-based explainability techniques, including SmoothGrad [48], GradInput [46], and GradCAM [44], 2) Bert with the above feature-based explainability techniques, where Group SELFIES is used as input for a fair comparison, and 3) explainable GNNs including OrphicX [30], GNNExplainer [64], PGExplainer [33], and Gem [29]. The details of parameter settings can be found in the supplemental material.
TABLE II: Mean Classification Accuracy on the Seven Datasets (%)
Methods Mutag Mutagen PTC-FM PTC-FR PTC-$M$ PTC-MR Liver
GCN [23] 84.6 78.9 54.8 63.0 57.8 53.3 41.1
DGCNN [66] 85.8 74.8 57.3 63.5 61.0 58.6 44.6
edGNN [21] 86.9 75.2 59.8 65.7 64.4 56.3 44.5
GIN [62] 87.5 82.3 62.1 66.2 65.1 64.0 44.9
RW-GNN [36] 87.2 80.3 61.9 64.0 62.4 57.0 43.2
DropGNN [37] 89.4 80.7 62.0 66.0 63.7 64.2 45.0
IEGN [34] 84.6 80.1 60.8 59.8 61.1 59.5 45.3
ChemBERTa [7] 86.8 78.0 60.0 65.7 60.4 58.7 45.7
@@
TABLE III: Mean Explanation Accuracy on the Seven Datasets (%)
Methods Mutag Mutagen PTC-FM PTC-FR PTC-$M$ PTC-MR Liver
GradInput+GCN [46] 70.3 67.9 69.7 66.4 64.6 65.0 73.0
GradCAM+GCN [44] 69.8 67.0 71.0 67.9 66.2 67.3 69.4
SmoothGrad+GCN [48] 69.2 66.8 67.5 62.6 64.9 63.1 66.4
GradInput+Bert [46] 75.1 72.6 73.0 68.9 65.6 69.6 75.6
GradCAM+Bert [44] 75.3 72.4 77.5 70.0 70.2 73.0 76.0
SmoothGrad+Bert [48] 73.4 72.8 73.7 71.0 67.0 69.9 75.1
GNNExplainer [64] 70.6 64.2 68.9 67.9 66.8 67.1 72.1
PGExplainer [33] 66.5 58.7 70.3 68.0 65.9 67.0 71.5
Gem [29] 73.7 66.0 71.3 69.0 68.9 69.2 73.6
OrphicX [30] 78.0 71.5 74.6 70.4 70.9 71.4 74.0
@@
TABLE IV: MEAN EXPLANATIONS’ FIDELITY ON THE SEVEN DATASETS
Methods Mutag Mutagen PTC-FM PTC-FR PTC-$M$ PTC-MR Liver
SmoothGrad+GCN [48] 0.31 0.28 0.30 0.25 0.29 0.27 0.33
CAM+GCN [67] 0.30 0.27 0.29 0.26 0.28 0.32 0.35
GradInput+GCN [46] 0.32 0.26 0.31 0.28 0.25 0.26 0.34
GradCAM+GCN [44] 0.33 0.29 0.34 0.30 0.31 0.33 0.36
SmoothGrad+Bert [48] 0.40 0.38 0.42 0.39 0.35 0.38 0.45
CAM+Bert [67] 0.39 0.37 0.36 0.38 0.34 0.37 0.43
GradInput+Bert [46] 0.42 0.40 0.41 0.36 0.35 0.39 0.46
GradCAM+Bert [44] 0.43 0.41 0.47 0.40 0.42 0.44 0.48
OrphicX [30] 0.45 0.43 0.44 0.40 0.41 0.43 0.47
GNNExplainer [64] 0.35 0.31 0.34 0.32 0.33 0.35 0.40
PGExplainer [33] 0.33 0.29 0.35 0.31 0.32 0.34 0.39
Gem [29] 0.41 0.36 0.38 0.37 0.39 0.40 0.44
Lamole+DistilBert 0.50 0.52 0.55 0.51 0.53 0.62 0.58
Lamole+DeBerta 0.58 0.60 0.65 0.55 0.56 0.61 0.59
Lamole+Bert 0.61 0.62 0.68 0.57 0.58 0.57 0.63
IV-B Results and Discussion
Prediction Performance. Table II shows the classification accuracy of the compared algorithms. As we can see, our proposed Lamole+DistilBert, Lamole+DeBerta, and Lamole+Bert not only can provide explainability but also can achieve comparable prediction accuracy as compared to existing predictive methods. In addition, Lamole models show superior performance over ChemBERTa. This suggests that using molecular representations with more chemical semantics, like Group SELFIES, can help LMs better learn the chemical semantics and structure-property relationships.
Explanation Performance. Table III presents the explanation accuracy of the compared explainability techniques. It should be noted that the ground truth annotations used in our work provide additional supervisory signals. Therefore, we also align the generated explanations of these baselines with the annotations for fair comparison, and the ground truth annotation rate is 10%. Lamole improves the explanation accuracy by $1.4\%\sim 14.3\%$ compared to the baseline methods.
The results of the explanation accuracy of Lamole under different annotation rates (10%, 20%, 50%, and 100%) on the PTC-FR and PTC-MM datasets are shown in Fig. 5 and Fig. 6 (a), (b). It is clear that more annotations can constantly enhance the accuracy of the explanation. Compared to Lamole ($\mathcal{L}_{M}$), only using 10% molecules with ground truth annotations can significantly improve explanation accuracy by up to 5%. Using more annotations (from 10% to 20%) can achieve significant improvement in explanation accuracy. However, raising the rate from 50% to 100% can bring a limited increase in explanation accuracy on the four PTC datasets. Therefore, there is a trade-off between the explanation accuracy and additional annotation costs.
The fidelity results across the seven datasets presented in Table IV confirm that our method, Lamole, generates explanations that are more faithful to the underlying model prediction than all baselines. The best performing Lamole+Bert, consistently achieved the highest scores on nearly all benchmarks, including 0.61 on Mutag, 0.62 on Mutagen, and 0.68 on PTC-FM. This performance is substantially higher compared to dedicated explainers like OrphicX and gradient methods. The data indicate that Lamole’s integration of attention and gradient signals, combined with its functional-group focus, produces explanations where the removal of the identified fragments causes a maximal reduction in the model’s prediction score. This confirms the reliability of our approach.
(a) Mutag
(b) PTC-FM
(c) PTC-MR
Figure 5: The explanation accuracy of Lamole with different annotation rates on the Mutag, PTC-FM, and PTC-MR datasets.
(a) PTC-FR
(b) PTC-$M$
(c) Molecule (ID: 155)
@@ IV-B Results and Discussion
Figure 6: The explanation accuracy of Lamole with different annotation rates on the PTC-FR and PTC-MM datasets, and the explanation visualization of one molecule from the Mutag dataset.
In addition, we discuss the rationale of using labeled annotations. Lamole requires human-labeled annotations. Due to the huge knowledge base of LLMs, we explored the use of LLMs, including ChatGPT and ChemLLM [65], to annotate the ground truth. We input molecules’ SMILES strings into the two LLMs to ask the ground truth. The explanation accuracy results are shown in Table V.
TABLE V: The explanation accuracy results when using different annotation methods
Methods Mutag PTC-FM PTC-FR PTC-$M$ PTC-MR Liver
Lamole (ChatGPT) 67.6 64.5 55.0 51.0 65.6 72.7
Lamole (ChemLLM) 62.5 72.1 61.5 57.1 71.9 72.7
Lamole (Human) 77.8 81.1 72.2 72.0 73.1 77.3
It is obvious that there is a significant decrease in explanation accuracy, indicating that existing LLMs may make incorrect annotations. Therefore, we argue that this human-in-the-loop strategy — providing slight human annotations — to guide learning is reasonable and necessary for critical scientific domains.
We selected some representative molecules for explanation visualization. These explanations are shown in Figs. 1, 6 (c), 7, and Fig. S-1 (see supplemental material), respectively. The right panel of those figures is the importance scores obtained by Lamole, where "other" in the figures is the average importance score of other unlisted functional groups/fragments. Compared to baseline methods, Lamole provides chemically meaningful explanations. Particularly, the interaction among the functional groups is successfully captured. As shown in Figs. 7 (b) and S-1 (a), Lamole accurately and confidently identify benzene with amido group and benzene-1 with nitro-1 group as explanations, respectively. While other methods can neither provide chemically meaningful explanations nor reflect the functional group interactions. These explanations demonstrate Lamole’s superior interpretation in faithfully revealing the structure-property relationships. However, as depicted in Fig. S-1 (b), although the ground truth substructures, i.e., benzene-1 and amido ground, are identified, other functional groups/fragments such as carbonyl-0, $\mathrm{Br}$-0, and $\mathrm{Br}$-1 also have relatively higher importance scores. This may be due to complex interactions caused by multiple functional groups. In future work, more strategies may need to be designed to reveal such complex functional group interactions.
(a) Molecule (ID: 156) from the Mutag dataset
(b) Molecule (ID: 574) from the Liver dataset
Figure 7: Explanation visualization molecules from the Mutag and the Liver datasets.
In addition, we evaluated the performance of the compared algorithms by using the proposed explanation plausibility metric $\mathrm{EP}$. The statistical results of $\mathrm{EP}$ on the PTC-FR, PTC-MM, and Liver datasets are presented in Figs. 8 and S-2. From the figures, we can observe that the $\mathrm{EP}$ values of the comparison algorithm are slightly lower, which means that the algorithms cannot confidently reflect the relationships between structure and property. Compared to the comparison algorithm, the $\mathrm{EP}$ values of Lamole have increased by $2\%\sim 9\%$. The explanations’ plausibility $\mathrm{EP}(g)$ is defined as the ratio of the difference between the mean importance scores of ground truth and the mean importance scores of non-ground truth to the mean importance scores of non-ground truth. A larger ratio indicates that ground truth’s importance scores exceed non-ground truth’s. The high $\mathrm{EP}(g)$ values of Lamole indicate Lamole can improve the importance scores of ground truth and suppress the importance scores of non-ground truth, leading to higher confidence in matching the ground truth.
@@
IV-C Applicability to Broader Properties
The above experiments focus on toxicity-related properties where some ground-truth concepts are well-understood. To further investigate whether the explainable prediction ability applies to broader properties, we have performed additional experiments on two critical ADME properties: solubility and permeability.
Since large-scale human-annotated datasets for ADME properties are unavailable, we employ an automated, descriptor-driven approach to generate the ground-truth functional group labels required for evaluating explanation accuracy on the ZINC dataset. We leverage established physicochemical principles quantified by the RDKit cheminformatics package [28] to define chemically meaningful key fragments. For both solubility and permeability, the process hinges on quantifying local contributions to defining global descriptors, such as topological Polar Surface Area (tPSA) for polarity and LogP for hydrophobicity. We utilize RDKit to calculate the contribution of each atom to the relevant property descriptor. We determine the ground truth structures by aggregating these atomic contributions: the contributions of all atoms belonging to a specific Group SELFIES fragment are summed. The fragments whose summed contributions are among the highest are then designated as the ground truth structures. Fragments with the highest tPSA are labeled as ground truth structures  [^1] It should be noted that the purpose of using this RDKit-based annotation is to provide computational references, rather than the true chemical causality — this needs to be rigorously verified through wet experiments..
The results presented in Table VI and Table VII demonstrate that the advantages of the Lamole framework extend to these more diverse and complex pharmacological tasks, confirming its applicability beyond toxicity prediction. Regarding predictive performance (Table VI), Lamole shows substantial superiority over GNN baselines in terms of Root Mean Square Error (RMSE). For instance, Lamole+Bert achieved an RMSE of 0.65 for solubility. A similar trend holds for permeability, where Lamole+Bert achieved an RMSE of 2.1, significantly outperforming GNN baselines. This predictive accuracy confirms that Lamole’s architecture captures the subtle structure-activity relationships crucial for accurate ADME prediction.
The results for explanation quality (Table VII) further show the method’s generalizability by yielding higher explanation accuracy and fidelity scores. The framework achieved the highest explanation accuracy (Exp. Acc.) for both solubility (89.1%) and permeability (73.5%), indicating that Lamole is superior at identifying the core functional groups that govern physicochemical behavior. Furthermore, the fidelity scores—which measure how faithful the explanation is to the model’s prediction—are also highest for Lamole variants. The explanation visualization shown in Fig. S-3 and the explanation plausibility shown in Fig. S-4 also confirmed the accuracy of the derived explanations.
TABLE VI: Comparison of Prediction Performance (RMSE) on Solubility and Permeability
Methods Solubility Permeability
GCN [23] 0.88 3.5
DGCNN [66] 0.85 3.4
edGNN [21] 0.82 3.3
GIN [62] 0.79 3.1
RW-GNN [36] 0.77 3.0
DropGNN [37] 0.75 2.9
IEGN [34] 0.73 2.9
Lamole+DistilBert 0.70 2.8
Lamole+DeBerta 0.74 2.5
Lamole+Bert 0.65 2.1
TABLE VII: Comparison of Explanation Quality on Solubility and Permeability
Methods Solubility Permeability
Exp. Acc. ($\%$) Fidelity Exp. Acc. ($\%$) Fidelity
SmoothGrad+GCN [48] 78.5 0.35 64.1 12.3
CAM+GCN [67] 79.1 0.38 65.0 13.5
GradInput+GCN [46] 77.8 0.32 63.8 11.9
GradCAM+GCN [44] 80.3 0.41 66.2 14.0
SmoothGrad+Bert [48] 81.5 0.43 67.5 14.5
CAM+Bert [67] 82.2 0.45 68.1 14.8
GradInput+Bert [46] 81.0 0.42 66.9 14.1
GradCAM+Bert [44] 83.0 0.46 69.5 15.1
OrphicX [30] 84.5 0.48 70.0 15.5
GNNExplainer [64] 85.1 0.49 70.8 15.8
PGExplainer [33] 83.9 0.47 69.8 15.3
Gem [29] 85.5 0.49 71.0 16.0
Lamole+DistilBert 88.4 0.52 72.1 16.5
Lamole+DeBerta 88.7 0.50 73.2 15.9
Lamole+Bert 89.1 0.56 73.5 17.6
IV-D Ablation Studies
[…]
Figure 9: The explanation accuracy of Lamole+Bert, Lamole+Bert ($\alpha$), Lamole+Bert ($\mathcal{L}_{M}$), and Lamole (Att).
[…]